OpenStudy (anonymous):

The molecular mass of nitrogen is 14 times greater than that of hydrogen. If the molecules in these two gases have the same rms speed, the ratio of hydrogens absolute temperature to that of nitrogen is?

OpenStudy (anonymous):

\[\sqrt{3RT _{1}/M}=\sqrt{3RT _{2}/14M}\] \[14M/M=T _{2}/T _{1}\] \[14T _{1}=T _{2}\] Therefore, temperature ratio of hydrogen to nitrogen is 14:1, meaning the hydrogen is 14 times the absolute temperature of nitrogen. I think